OpenStudy (thickgirl_angel):

Zeus Industries bought a computer for $2857. It is expected to depreciate at a rate of 24% per year. What will the value of the computer be in 3 years?

OpenStudy (welshfella):

start again!! Value after x years = 2487 - 2487*0.24x Plug x = 3 into the above to find the value after 3 years.

OpenStudy (welshfella):

{note 24 % = 0.24 }

OpenStudy (welshfella):

so just replace the x in the formula by 3 and work it out on your calculator.

OpenStudy (welshfella):

can you do that?

OpenStudy (welshfella):

note that 0.24x means ' 0.24 times x' so here you would use '0.24 times 3'.

OpenStudy (thickgirl_angel):

0.72?

OpenStudy (welshfella):

yea so putting that into the full equation we have value after 3 years = 2487 - 2487*0.72